OpenStudy (anonymous):

v=√2.5r The equation describes the relationship between the radius r in feet of an unbanked curve and the maximum velocity v in miles per hour that a car can safely go around the curve. A primary road with a speed limit of 35 miles per hour curves through a residential area. What is the radius of the unbanked curve? a) 490 ft b) 14 ft c) 49 ft d) 450 ft

OpenStudy (anonymous):

.v=35 v=sqrt(2.5r) 35=sqrt(2.5r)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Wait aha! Are those the multiple choices?

OpenStudy (anonymous):

Yess

OpenStudy (ankit042):

The max speed is 35 miles/hr and and v= sqrt(2.5)r describnes the radius and speed relationship so you calculate r first but it will be in miles so convert to get it in feets

OpenStudy (anonymous):

I think its C not sure thou